a pizza with a 25cm diameter cost $4.60.
a pizza with a 31cm diameter cost $6.00.
How do you work how which has the best value?

Found 2 solutions by mananth, rfer:
Answer by mananth(16946)

25cm diameter cost $4.60.
area of pizza = pi*r^2
diameter = 25 cm . radius = 12.5 cm
A = pi*(12.5)^2
= 491 sq. cm
cost = $4.60
cost per unit area = 4.6/491
= 0.93 cents / sq.cm
..
a pizza with a 31cm diameter cost $6.00.
A= pi*(15.5)^2
= 754.86 sq. cm
cost per sq.cm = 6.00/754.86
0.79 cents / sq. cm.----------better value

Answer by rfer(16322)

A=3.14r^2
A=3.14*156.25
A=490.625 sq cm/4.60=106.67 sq cm per dollar
-------
A=3.14*240.25
A=754.385 sq cm/6=125.73 sq cm per dollar
----------
Find the unit price to compare.